# T-Shirt: Token-Selective Hierarchical Data Selection for Instruction Tuning

### v1

@@ Abstract
Instruction tuning is essential for Large Language Models (LLMs) to effectively follow user instructions. To improve training efficiency and reduce data redundancy, recent works use LLM-based scoring functions, e.g., Instruction-Following Difficulty (IFD), to select high-quality instruction-tuning data with scores above a threshold. While these data selection methods often lead to models that can match or even exceed the performance of models trained on the full datasets, we identify two key limitations: (i) they assess quality at the sample level, ignoring token-level informativeness; and (ii) they overlook the robustness of the scoring method, often selecting a sample due to superficial lexical features instead of its true quality. In this work, we propose Token-Selective HIeRarchical Data Selection for Instruction Tuning (T-Shirt), a novel data selection framework that introduces a new scoring method to include only informative tokens in quality evaluation and also promotes robust and reliable samples whose neighbors also show high quality with less local inconsistencies. We demonstrate that models instruction-tuned on a curated dataset (only 5% of the original size) using T-Shirt can outperform those trained on the entire large-scale dataset by up to 5.48 points on average across eight benchmarks. Across various LLMs and training set scales, our method consistently surpasses existing state-of-the-art data selection techniques, while also remaining both cost-effective and highly efficient. For instance, by using GPT-2 for score computation, we are able to process a dataset of 52k samples using 40 minutes on a single GPU.
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable success in various tasks [1, 2, 3, 4, 5]. One key component of adapting these powerful LLMs for practical use is instruction tuning [6] that fine-tunes pre-trained LLMs on instruction–response pairs. During this stage, researchers often use millions of samples [3, 4, 7]. However, LIMA [8] challenges this norm with the Superficial Alignment Hypothesis, which indicates that instruction tuning might not need to be data-intensive, shifting the focus from data quantity to data quality. Fine-tuned on just 1,000 manually selected instruction-tuning samples, LIMA achieves strong performance. Data selection for instruction tuning offers several key benefits. First, full-dataset tuning is extremely time-consuming and costly. Furthermore, as synthesized data becomes more common [9, 10, 11, 12], datasets often include redundancy and noise, making data selection for instruction tuning absolutely essential.
@@
To select instruction tuning data, it is common to use scoring functions to assess sample quality and only keep samples with high scores. LIMA [8] relies on manual annotation, while others [13, 14, 15] prompt proprietary LLMs via APIs to assign quality scores. However, both methods are costly. An alternative is Instruction-Following Difficulty (IFD) score [16, 17] (elaborated in Equation 1), which measures the ratio of a response’s perplexity when conditioned on the instruction to its unconditioned perplexity. Essentially, higher IFD scores indicate better quality. IFD is quite appealing due to its efficiency and low cost, and even small models like GPT-2 [18] can be used to compute it.
@@
However, current scoring functions have multiple limitations as we identify in this work. First, they evaluate the training data at the sample level, overlooking finer-grained token-level information. Recent studies show that not all tokens contribute equally during training [19, 20]. Instruction tuning significantly affects models’ output on only a small subset of response tokens [21, 22]. These findings motivate us to reconsider previous sample-level scoring paradigms. We revisit the computation of IFD scores to examine the informativeness of individual tokens. We find that the IFD score treats all response tokens equally, and a sample where none of the individual tokens are informative for instruction tuning may still receive a high overall score (see our counterexample in Figure 2). Empirically, we find that over 20% of the response tokens in instruction tuning datasets are not informative. Excluding these tokens from IFD calculation can substantially impact the final score.
@@
Second, existing works using LLM-based quality scores overlook the robustness of quality assessment, particularly with respect to the local neighborhood of a training sample. These methods typically apply a fixed threshold and treat samples with scores above it as high quality. However, fluctuations may exist within a sample’s neighborhood, where some of its neighboring samples fall below the threshold. We find that the IFD score is not a robust metric. Small, semantic-preserving changes, such as replacing a word with its synonym, can lead to large shifts in the IFD score (see our counterexample in Figure 4 with more details in Section 3.2). This suggests that high scores may also result from reliance on spurious features rather than reflecting true sample quality, motivating the need to also examine the local neighborhood of a sample during quality assessment.
Figure 1: An overview of our approach, T-Shirt.$W_{E}$ denotes the model’s embedding layer (e.g., GPT-2) used to compute S-IFD scores. For each instruction-response pair, we generate its neighbors by perturbing token embeddings. Then we only use selected response tokens (green squares) for S-IFD computation, while excluded tokens are marked in red squares. Finally, we use hierarchical selection to choose samples whose neighbors exhibit high average S-IFD and low variance.
@@
To address the two challenges above, we propose T-Shirt, short for Token-Selective HIeRarchical Data Selection for Instruction Tuning, which is a new framework for data selection that promotes both token-level informativeness and robust quality assessment (also see Figure 1 for an overview). Towards introducing our framework, we first propose Selective IFD (S-IFD) (Equation 3) that evaluates each response token individually and includes only informative tokens in the final quality score. Next, to avoid selecting samples whose high scores result from superficial lexical cues rather than true semantic quality, we go beyond previous pipelines that naively select samples based on a fixed quality score threshold. Instead, we apply a hierarchical selection strategy that favors samples whose neighbors exhibit high average and low variance in S-IFD scores. Experiments across multiple instruction tuning datasets and pretrained LLMs show that our proposed method T-Shirt outperforms existing baselines on a wide range of downstream tasks. At the same time, our method remains computationally and financially efficient. For example, T-Shirt requires only about 40 minutes to select data from the 52k-sample Alpaca-GPT-4 [11] dataset using GPT-2 on a single GPU.
@@ 1.1 Related Work
Data Selection for Instruction TuningLIMA [8] challenges the need for data-intensive instruction tuning [3, 4, 7] by proposing the Superficial Alignment Hypothesis (SAH) and showing that a 65B model instruction-tuned on just 1,000 manually curated samples can outperform models trained with massive data. To reduce the cost of human annotation, subsequent research explored automated quality scoring functions for data selection. Early methods use simple metrics such as response length [23] and perplexity [24]. Later approaches rely on proprietary LLMs for quality scoring. For example, AlpaGasus [13] and DS2 [15] use OpenAI APIs, while Deita [14] trains custom LLM scorers using ChatGPT-generated labels. To avoid expensive API calls, researchers propose using the IFD score [16], which can be computed even with small models like GPT-2 [17]. Other methods assess data quality via gradient-based influence functions [25] and Shapley values [26], but these approaches are computation-heavy and require additional validation sets. Besides sample quality, some works like LIMA [8], Deita [14] and DS2 [15], also emphasize sample diversity. However, most existing methods evaluate data quality at the sample level, overlooking token-level quality. They also do not consider the robustness of quality scores to small input variations.
@@
Token-level Informativeness  Recent works [21, 22] further support SAH by showing that alignment primarily affects models’ outputs on a small subset of response tokens. This leads us to hypothesize that token-level analysis might be important for data selection. Other studies [19, 20] show that not all tokens contribute equally during pre-training and instruction tuning. Then they use selective language modeling (SLM) to alter the training process itself by computing losses only on selected tokens. Selecting tokens requires training a reference model on high-quality data and comparing token losses between the untrained and reference models. This makes data selection a prerequisite for building the reference model. In contrast, our metric, S-IFD, does not need a reference model and works well even if it is computed by GPT-2. Moreover, our method operates at the data preparation stage and does not change the training process. Thus, our method is orthogonal to SLM, and it has the potential to be combined with SLM for further improvement on instruction tuning.
@@
Robustness for LLMs  Theoretically, LLMs are not robust to minor input perturbations because the self-attention mechanism [27] is not smooth and Lipschitz continuous [28]. Empirically, even small and semantic-preserving input changes can lead LLMs to misclassify text [29, 30, 31] or generate incorrect code [32]. Currently, robustness in LLM-based scoring functions for data selection remains underexplored. To the best of our knowledge, we are the first to incorporate robustness assessment into data selection for instruction tuning.
@@ 2 Preliminaries
Scoring Function  Selection policies typically use scoring functions to evaluate each instruction-response pair in the dataset $\mathcal{D}$, favoring samples with higher scores. Some functions focus on a single aspect, such as quality or diversity of samples [13, 16, 17, 23], while others assess multiple attributes [8, 14, 15]. In this work, we focus solely on quality, using a well-established metric called Instruction-Following Difficulty (IFD) [16]. It is a desirable metric because it is purely statistical and easy to compute. For a given instruction-response pair $(x,y)$, the IFD score is computed using an LLM with parameters $\theta^{\prime}$. It is defined as the ratio of the perplexity ($\text{PPL}_{\theta^{\prime}}$) of the response conditioned on the instruction to the unconditional perplexity of the response:
@@
where $T$ is the length of the response, $y_{t}$ is the $t$-th token in the response, $y_{<t}$ represents all response tokens before the $t$-th token, and $P_{\theta^{\prime}}$ is the probability assigned by the model $\theta^{\prime}$ to the token $y_{t}$. A higher value of $P_{\theta^{\prime}}$ indicates that the token $y_{t}$ is easy to predict under the model. The IFD score measures how much the instruction $x$ helps the model generate the response $y$. A lower score indicates that the instruction made the model easier to generate the response $y$, while a higher score suggests greater difficulty and thus a more informative sample for fine-tuning. However, samples with $\text{IFD}(x,y)\geq 1$ are discarded, as they indicate that the instruction does not aid the response generation at all. The model used to compute IFD scores can be the same as the model being instruction-tuned ($\theta^{\prime}=\theta$) or a different one ($\theta^{\prime}\neq\theta$). Superfiltering [17] shows that even small models like GPT-2 can be used for computing IFDs.
@@ 3.1 Selective IFD to Capture Token-Level Informativeness
[…]
LLMs acquire most of their knowledge during pretraining. So even without the instruction, the token following a response prefix can be easy for them to predict. As shown in Figure 2(a), when tokens “famous scientist” appear early in the response, pretrained LLMs can easily predict that the next token after “Albert” is “Einstein”, regardless of whether the instruction is “Identify a famous scientist.” or “Who proposed the theory of relativity?” This suggests that not all tokens are informative for instruction following. Therefore, we reconsider current data selection approaches, which typically evaluate instruction-response pairs as whole units without examining each token individually, and propose the question: “Within an instruction-response pair, how informative is each response token?
To answer this, we revisit the IFD score. Using $\Delta_{t}$ to denote $\log P_{\theta^{\prime}}\left(y_{t}|y_{<t},x\right)-\log P_{\theta^{\prime}}\left(y_{t}|y_{<t}\right)$, we first rewrite Equation 1 as follows:
@@
The IFD score treats all tokens equally, which might be misleading. We use the following thought experiment to demonstrate the potential issue intuitively. If for all $y_{t}$ in an instruction-response pair, $\Delta_{t}\approx 0.01$, the IFD score would be approximately 0.99, suggesting that this training sample has very high quality. However, at the token level, none of these tokens significantly depend on the instruction, indicating low informativeness. Empirically, we find that a large fraction of response tokens in instruction tuning datasets exhibit small $|\Delta_{t}|$ values, regardless of the choice of $\theta^{\prime}$ for computing IFD scores. For instance, in the Alpaca-GPT-4 dataset [11], using GPT-2 as $\theta^{\prime}$, about 22% of response tokens have $|\Delta_{t}|\leq 0.01$, and this proportion is around 28% for both Llama-3.1-8B [3] and Qwen-2.5-7B [4]. We further illustrate this by selecting two samples from the Alpaca-GPT-4 dataset and computing token-level $|\Delta_{t}|$ using GPT-2. As shown at the bottom of Figure 2, we observe that $|\Delta_{t}|$ for most tokens is small.
@@
where $\text{where }w_{t}=\begin{cases}1&\text{ if }|\Delta_{t}|\text{ ranks top }k\%\text{ in the dataset }\mathcal{D},\\ 0&\text{ otherwise}.\end{cases}$
We illustrate the distinction between IFD and S-IFD using two examples in Figure 2, as they share similar IFD scores but exhibit notably different S-IFD values. We view that the second example, with a higher S-IFD score, has better quality.

### v2

@@ Abstract
Instruction tuning is essential for Large Language Models (LLMs) to effectively follow user instructions. To improve training efficiency and reduce data redundancy, recent works use LLM-based scoring functions, e.g., Instruction-Following Difficulty (IFD), to select high-quality instruction-tuning data with scores above a threshold. While these data selection methods often lead to models that can match or even exceed the performance of models trained on the full datasets, we identify two key limitations: (i) they assess quality at the sample level, ignoring token-level informativeness; and (ii) they overlook the robustness of the scoring method, often selecting a sample due to superficial lexical features instead of its true quality. In this work, we propose Token-Selective HIeRarchical Data Selection for Instruction Tuning (T-Shirt), a novel data selection framework that introduces a new scoring method to include only informative tokens in quality evaluation and also promotes robust and reliable samples whose neighbors also show high quality with less local inconsistencies. We demonstrate that models instruction-tuned on a curated dataset (only 5% of the original size) using T-Shirt can outperform those trained on the entire large-scale dataset by up to 5.48 points on average across eight benchmarks. Across various LLMs and training set scales, our method consistently surpasses existing state-of-the-art data selection techniques, while also remaining both cost-effective and highly efficient. For instance, by using GPT-2 for score computation, we are able to process a dataset of 52k samples in 40 minutes on a single GPU. Our code is available at https://github.com/Dynamite321/T-SHIRT.
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable success in various tasks [brown2020gpt3, openai2024gpt4technicalreport, grattafiori2024llama3herdmodels, qwen2025qwen25technicalreport, deepseek2005deepseekr1]. One key component of adapting these powerful LLMs for practical use is instruction tuning (also known as supervised fine-tuning, SFT) [wei2022finetuned] that fine-tunes pre-trained LLMs on instruction–response pairs. During this stage, researchers often use millions of samples [grattafiori2024llama3herdmodels, qwen2025qwen25technicalreport, lambert2025tulu3]. However, LIMA [zhou2023lima] challenges this norm with the Superficial Alignment Hypothesis, which indicates that instruction tuning might not need to be data-intensive, shifting the focus from data quantity to data quality. Fine-tuned on just 1,000 manually selected instruction-tuning samples, LIMA achieves strong performance. Data selection for instruction tuning offers several key benefits. First, full-dataset tuning is extremely time-consuming and costly. Furthermore, as synthesized data becomes more common [wang2023selfinstruct, xu2025magpie, peng2023instructiontuninggpt4, xu2024wizardlm], datasets often include redundancy and noise, making data selection for instruction tuning absolutely essential.
@@
To select instruction tuning data, it is common to use scoring functions to assess sample quality and only keep samples with high scores. LIMA [zhou2023lima] relies on manual annotation, while others [chen2024alpagasus, liu2024deita, pang2025ds2] prompt proprietary LLMs via APIs to assign quality scores. However, both methods are costly. An alternative is Instruction-Following Difficulty (IFD) score [li2024ifd, li2024superfiltering] (elaborated in Equation˜1), which measures the ratio of a response’s perplexity when conditioned on the instruction to its unconditioned perplexity. Essentially, higher IFD scores indicate better quality. IFD is quite appealing due to its efficiency and low cost, and even small models like GPT-2 [radford2019gpt2] can be used to compute it.
@@
However, current scoring functions have multiple limitations as we identify in this work. First, they evaluate the training data at the sample level, overlooking finer-grained token-level information. Recent studies show that not all tokens contribute equally during training [lin2024rho1, pang2025tokencleaning]. Instruction tuning significantly affects models’ output on only a small subset of response tokens [qi2025safety, lin2024unlocking]. These findings motivate us to reconsider previous sample-level scoring paradigms. We revisit the computation of IFD scores to examine the informativeness of individual tokens. We find that the IFD score treats all response tokens equally, and a sample where none of the individual tokens are informative for instruction tuning may still receive a high overall score (see our counterexample in Figure˜2). Empirically, we find that over 20% of the response tokens in instruction tuning datasets are not informative. Excluding these tokens from IFD calculation can substantially impact the final score.
@@
Second, existing works using LLM-based quality scores overlook the robustness of quality assessment, particularly with respect to the local neighborhood of a training sample. These methods typically apply a fixed threshold and treat samples with scores above it as high quality. However, fluctuations may exist within a sample’s neighborhood, where some of its neighboring samples fall below the threshold. We find that the IFD score is not a robust metric. Small, semantic-preserving changes, such as replacing a word with its synonym, can lead to large shifts in the IFD score (see our counterexample in Figure˜4 with more details in Section˜3.2). This suggests that high scores may also result from reliance on spurious features rather than reflecting true sample quality, motivating the need to also examine the local neighborhood of a sample during quality assessment.
[…]
@@
To address the two challenges above, we propose T-Shirt, short for Token-Selective HIeRarchical Data Selection for Instruction Tuning, which is a new framework for data selection that promotes both token-level informativeness and robust quality assessment (also see Figure˜1 for an overview). Towards introducing our framework, we first propose Selective IFD (S-IFD) (Equation˜3) that evaluates each response token individually and includes only informative tokens in the final quality score. Next, to avoid selecting samples whose high scores result from superficial lexical cues rather than true semantic quality, we go beyond previous pipelines that naively select samples based on a fixed quality score threshold. Instead, we apply a hierarchical selection strategy that favors samples whose neighbors exhibit high average and low variance in S-IFD scores. Experiments across multiple instruction tuning datasets and pretrained LLMs show that our proposed method T-Shirt outperforms existing baselines on a wide range of downstream tasks. At the same time, our method remains computationally and financially efficient. For example, T-Shirt requires only about 40 minutes to select data from the 52k-sample Alpaca-GPT-4 [peng2023instructiontuninggpt4] dataset using GPT-2 on a single GPU.
@@ 1.1 Related Work
Data Selection for Instruction TuningLIMA [zhou2023lima] challenges the need for data-intensive instruction tuning [grattafiori2024llama3herdmodels, qwen2025qwen25technicalreport, lambert2025tulu3] by proposing the Superficial Alignment Hypothesis (SAH) and showing that a 65B model instruction-tuned on just 1,000 manually curated samples can outperform models trained with massive data. To reduce the cost of human annotation, subsequent research explored automated quality scoring functions for data selection. Early methods use simple metrics such as response length [zhao2024long] and perplexity [cao2024instructionmining]. Later approaches rely on proprietary LLMs for quality scoring. For example, AlpaGasus [chen2024alpagasus] and DS2 [pang2025ds2] use OpenAI APIs, while Deita [liu2024deita] trains custom LLM scorers using ChatGPT-generated labels. To avoid expensive API calls, researchers propose using the IFD score [li2024ifd], which can be computed even with small models like GPT-2 [li2024superfiltering]. Other methods assess data quality via gradient-based influence functions [xia2024less] and Shapley values [he2024shed], but these approaches are computation-heavy and require additional validation sets. Besides sample quality, some works like LIMA [zhou2023lima], Deita [liu2024deita] and DS2 [pang2025ds2], also emphasize sample diversity. However, most existing methods evaluate data quality at the sample level, overlooking token-level quality. They also do not consider the robustness of quality scores to small input variations.
@@
Token-level Informativeness  Recent works [qi2025safety, lin2024unlocking] further support SAH by showing that alignment primarily affects models’ outputs on a small subset of response tokens. This leads us to hypothesize that token-level analysis might be important for data selection. Other studies [lin2024rho1, pang2025tokencleaning] show that not all tokens contribute equally during pre-training and instruction tuning. Then they use selective language modeling (SLM) to alter the training process itself by computing losses only on selected tokens. Selecting tokens requires training a reference model on high-quality data and comparing token losses between the untrained and reference models. This makes data selection a prerequisite for building the reference model. In contrast, our metric, S-IFD, does not need a reference model and works well even if it is computed by GPT-2. Moreover, our method operates at the data preparation stage and does not change the training process. Thus, our method is orthogonal to SLM, and it has the potential to be combined with SLM for further improvement on instruction tuning.
@@
Robustness for LLMs  Theoretically, LLMs are not robust to minor input perturbations because the self-attention mechanism [Vaswani2017attention] is not smooth and Lipschitz continuous [kim2021lipschitz, hamman2025quantifying]. Empirically, even small and semantic-preserving input changes can lead LLMs to misclassify text [jin2020isbert, zang2020word, wang2021adversarial] or generate incorrect code [wang2023recode]. Currently, robustness in LLM-based scoring functions for data selection remains underexplored. To the best of our knowledge, we are the first to incorporate robustness assessment into data selection for instruction tuning.
@@ 2 Preliminaries
Scoring Function  Selection policies typically use scoring functions to evaluate each instruction-response pair in the dataset $\mathcal{D}$, favoring samples with higher scores. Some functions focus on a single aspect, such as quality or diversity of samples [chen2024alpagasus, li2024ifd, li2024superfiltering, zhao2024long], while others assess multiple attributes [zhou2023lima, liu2024deita, pang2025ds2]. In this work, we focus solely on quality, using a well-established metric called Instruction-Following Difficulty (IFD) [li2024ifd]. It is a desirable metric because it is purely statistical and easy to compute. For a given instruction-response pair $(x,y)$, the IFD score is computed using an LLM with parameters $\theta^{\prime}$. It is defined as the ratio of the perplexity ($\text{PPL}_{\theta^{\prime}}$) of the response conditioned on the instruction to the unconditional perplexity of the response:
@@
where $T$ is the length of the response, $y_{t}$ is the $t$-th token in the response, $y_{<t}$ represents all response tokens before the $t$-th token, and $P_{\theta^{\prime}}$ is the probability assigned by the model $\theta^{\prime}$ to the token $y_{t}$. A higher value of $P_{\theta^{\prime}}$ indicates that the token $y_{t}$ is easy to predict under the model. The IFD score measures how much the instruction $x$ helps the model generate the response $y$. A lower score indicates that the instruction made the model easier to generate the response $y$, while a higher score suggests greater difficulty and thus a more informative sample for fine-tuning. However, samples with $\text{IFD}(x,y)\geq 1$ are discarded, as they indicate that the instruction does not aid the response generation at all. The model used to compute IFD scores can be the same as the model being instruction-tuned ($\theta^{\prime}=\theta$) or a different one ($\theta^{\prime}\neq\theta$). Superfiltering [li2024superfiltering] shows that even small models like GPT-2 can be used for computing IFDs.
@@ 3.1 Selective IFD to Capture Token-Level Informativeness
(a) Example 1: high IFD but low $\text{S-IFD}_{75}$.
(b) Example 2: high IFD and high $\text{S-IFD}_{75}$.
Figure 2: Two examples from the Alpaca-GPT-4 dataset with nearly identical IFD scores but markedly different $\text{S-IFD}_{75}$ scores.Top: Instructions and partial responses from two examples, with tokens highlighted where $|\Delta_{t}|\leq 0.01$. Bottom: Plots of $|\Delta_{t}|$ values corresponding to the same examples. Tokens highlighted above are marked as red dots in the plots.
LLMs acquire most of their knowledge during pretraining. So even without the instruction, the token following a response prefix can be easy for them to predict. As shown in Figure˜2(a), when tokens “famous scientist” appear early in the response, pretrained LLMs can easily predict that the next token after “Albert” is “Einstein”, regardless of whether the instruction is “Identify a famous scientist.” or “Who proposed the theory of relativity?” This suggests that not all tokens are informative for instruction following. Therefore, we reconsider current data selection approaches, which typically evaluate instruction-response pairs as whole units without examining each token individually, and propose the question: “Within an instruction-response pair, how informative is each response token?
To answer this, we revisit the IFD score. Using $\Delta_{t}$ to denote $\log P_{\theta^{\prime}}\left(y_{t}|y_{<t},x\right)-\log P_{\theta^{\prime}}\left(y_{t}|y_{<t}\right)$, we first rewrite Equation˜1 as follows:
@@
The IFD score treats all tokens equally, which might be misleading. We use the following thought experiment to demonstrate the potential issue intuitively. If for all $y_{t}$ in an instruction-response pair, $\Delta_{t}\approx 0.01$, the IFD score would be approximately 0.99, suggesting that this training sample has very high quality. However, at the token level, none of these tokens significantly depend on the instruction, indicating low informativeness. Empirically, we find that a large fraction of response tokens in instruction tuning datasets exhibit small $|\Delta_{t}|$ values, regardless of the choice of $\theta^{\prime}$ for computing IFD scores. For instance, in the Alpaca-GPT-4 dataset [peng2023instructiontuninggpt4], using GPT-2 as $\theta^{\prime}$, approximately 22% of response tokens have $|\Delta_{t}|\leq 0.01$, as shown in Figure˜5 of Section˜C.1. This proportion is around 28% for both Llama-3.1-8B [grattafiori2024llama3herdmodels] and Qwen-2.5-7B [qwen2025qwen25technicalreport]. We further illustrate this by selecting two samples from the Alpaca-GPT-4 dataset and computing token-level $|\Delta_{t}|$ using GPT-2. As shown at the bottom of Figure˜2, we observe that $|\Delta_{t}|$ for most tokens is small.
@@
$\text{S-IFD}_{k}\left(x,y\right)=\exp\left\{-\frac{1}{\sum_{t=1}^{T}w_{t}}\sum_{t=1}^{T}w_{t}\Delta_{t}\right\},$ (3)
@@
We illustrate the distinction between IFD and S-IFD using two examples in Figure˜2, as they share similar IFD scores but exhibit notably different S-IFD values. We view that the second example, with a higher S-IFD score, has better quality.